the promblem is very confusing


The average annual undergraduated tution and fee charges C, in dollars for a certain 4 yr college for academic years 2000-2001 through 2005-2006 can be estmiated by the equation C=29.7x^2 +314.4x +3467.1, where x is the number of years after the 2000-2001 scademic year. assuming the model will remain valid beyond 2005-2006, in which academic YEAR will tution and fee charges be 8000?

Answer by stanbon(75887)

The average annual undergraduated tution and fee charges C, in dollars for a certain 4 yr college for academic years 2000-2001 through 2005-2006 can be estmiated by the equation C=29.7x^2 +314.4x +3467.1, where x is the number of years after the 2000-2001 scademic year. assuming the model will remain valid beyond 2005-2006, in which academic YEAR will tution and fee charges be 8000?
----------------------------
C(x) = 29.7x^2 +314.4x +3467.1
----
Solve 29.7x^2 + 314.4x + 3467.1 = 8000
---
29.7x^2 + 314.4x - 4532.9 = 0
----
I graphed the quadratic and found a positive
solution at x = 8.1472
---
Year = 2000+8.1472 means academic year 2009-10
====================================================
Cheers,
Stan H.
====================